# UmbraTTS: Adapting Text-to-Speech to Environmental Contexts with Flow Matching

### v1

@@ 2 Related Works
Recently, audio generation has advanced with models such as AudioLDM (Liu et al., 2023), AudioLDM2 (Liu et al., 2023), AudioGen (Kreuk et al., 2022), Make-An-Audio (Huang et al., 2023) and Fugatto (Valle et al., ). These models can create high-quality audio and soundscapes from text or other inputs. A specific subset of these models is environmentally-aware TTS models, which have mostly been explored through diffusion-based approaches, such as VoiceLDM (Lee et al., 2024), VoiceDiT (Jung et al., 2024) and Speak-in-the-Scene framework (Kim et al., 2024). Flow matching architectures (Lipman et al., 2022) have demonstrated superior synthesis quality in generative models, specifically for text-to-speech (TTS) tasks. Models like E2-TTS (Eskimez et al., 2024), F5-TTS (Chen et al., 2024), MatCha-TTS (Mehta et al., 2024), and Voiceox (Le et al., 2024) use direct probability paths for high-fidelity generation, but with no explicit environmental conditioning.
@@ 4.1 Quantitative Evaluation
First, we evaluated our model’s capability to generate speech with integrated environmental sounds. To assess this, we use the AudioCaps test subset as described in VoiceLDM (Lee et al., 2024), which contains audio recordings with both speech and background noise. During evaluation, our model is conditioned on the target text and the background sound from the test subset. The results are presented in Table 1. For quantitative evaluation, we report the Frechet Audio Distance (FAD) (Kilgour et al., 2018), Kullback-Leiber (KL), word error rate (WER), computed using Whisper large-v3 (Radford et al., 2023), and the CLAP score (Wu* et al., 2023), measured between the caption and the generated audio. The results show that UmbraTTS achieves SOTA performance compared to commonly used baselines. Next, we evaluated our model’s audio-to-audio capabilities. In this experiment, we conditioned our model only on the environmental target audio. The results are present in Table 2 for both the AudioCaps and MusicCaps. Again, our method achieves SOTA results across all evaluation metrics except for the FAD on AudioCaps, where it remains competitive.

### v2

@@ 2 Related Works
Recently, audio generation has advanced with models such as AudioLDM (Liu et al., 2023), AudioLDM2 (Liu et al., 2023), AudioGen (Kreuk et al., 2022), Make-An-Audio (Huang et al., 2023) and Fugatto (Valle et al., ). These models can create high-quality audio and soundscapes from text or other inputs. A specific subset of these models is environmentally-aware TTS models, which have mostly been explored through diffusion-based approaches, such as VoiceLDM (Lee et al., 2024), VoiceDiT (Jung et al., 2024) and Speak-in-the-Scene framework (Kim et al., 2024). Flow matching architectures (Lipman et al., 2022) have demonstrated superior synthesis quality in generative models, specifically for speech synthesis tasks (Chen et al., 2024; Navon et al., 2025). Models like E2-TTS (Eskimez et al., 2024), MatCha-TTS (Mehta et al., 2024), and Voiceox (Le et al., 2024) use direct probability paths for high-fidelity generation, but with no explicit environmental conditioning.
@@ 4.1 Quantitative Evaluation
First, we evaluated our model’s capability to generate speech with integrated environmental sounds. To assess this, we use the AudioCaps test subset as described in VoiceLDM (Lee et al., 2024), which contains audio recordings with both speech and background noise. During evaluation, our model is conditioned on the target text and the background sound from the test subset. The results are presented in Table 1. For quantitative evaluation, we report the Frechet Audio Distance (FAD) (Kilgour et al., 2018), Kullback-Leiber (KL), word error rate (WER), computed using Whisper large-v3 (Radford et al., 2023), and the CLAP score (Wu et al., 2023), measured between the caption and the generated audio. The results show that UmbraTTS achieves SOTA performance compared to commonly used baselines. Next, we evaluated our model’s audio-to-audio capabilities. In this experiment, we conditioned our model only on the environmental target audio. The results are present in Table 2 for both the AudioCaps and MusicCaps. Again, our method achieves SOTA results across all evaluation metrics except for the FAD on AudioCaps, where it remains competitive.